A dorm room at a college your sister is considering seems small. The room is rectangular and has a perimeter of 42 feet. The room is 4 feet lnger than it is wide. Ware the dimensions of the room?
Answer by scott8148(5896)

let x=width, so x+4=length __ p=2L+2W

42=2(x)+2(x+4) __ 42=2x+2x+8 __ 34=4x